# Overcoming Data Scarcity in Generative Language Modelling for Low-Resource Languages: A Systematic Review

### v1

@@ 4.1. RQ1: What Low-Resource Languages Have Already Been Modelled by Other Studies?
Many LRLs have been modelled by studies in this review for a variety of different natural language generation tasks. Given that Indo-European languages dominate the literature as the most frequently modelled LRLs with 133 different Indo-European languages being modelled, this could suggest that even among LRLs there is a disparity in terms of their representation. While many LRLs being modelled belong to the Indo-European language family, there are some language families such as the Aymaran and the Tupian language families whereby only a single language was modelled from their language family. It is worth considering that this disparity could be attributed to region-specific language families lacking the necessary resources to produce research in this area. This disparity indicates that there is a need for more research that is representative of all language families within the LRL group. Although this review is focused on bridging the data gap for modelling LRLs, there are other barriers such as research demand and sourcing expertise in the area. This could indicate that categorising the resources of a language requires considering the research interest, in addition to considering the amount of training data available for such an LRL.
@@
A wide range of technical methods have been used to overcome the challenge of data scarcity when building generative language models for LRLs. Similar to the results section for RQ2, the following discussion is divided by technical methods used to overcome data scarcity in order of frequency, starting with the most commonly used approaches.
@@ 4.2.1. Monolingual Text Data Augmentation
The process of synthesising new data from monolingual text data was the most frequently used approach for overcoming data scarcity. This often involved performing transformations on existing sentences. Some of these synthesis methods relate to replacing low-frequency terms with other terms chosen by an LSTM model (Fadaee et al., 2017; Haque et al., 2021; Shi et al., 2022). These approaches are bounded by the performance of the LSTM model chosen for replacement. Errors in the generated text produced by the chosen LSTM model could propagate to future models. In the context of language generation tasks such as question-answering and paraphrase generation, rare words could be indicative of domain-specific information or carry contextual information that would be lost with an LSTM replacement. This could contribute to nonsensical or vague information being produced by the models. Similarly, instead of performing a replacement on low-frequency terms, other studies have exploited POS tagging (Maimaiti et al., 2021) to inform the replacement process. While this method of replacing specific POS tags with similar terms using a paraphrase table could work for some LRLs, it could be challenging to apply this to LRLs with a complex morphological language such as Gaeilge, whereby altering a term in a sentence could corrupt the grammatical agreement of the sentence (Legate, 1999). This POS-informed replacement could be improved by adding a final neural approach for grammar correction. Moreover, another approach involves randomly replacing a word with a similar meaning using a pre-trained language model (Haque et al., 2021; Gao et al., 2019). This method could be enhanced by combining it with the previously mentioned POS-informed replacement. Performing contextualised replacements on specific POS tags could circumvent the issue of agreement and other grammatical errors. Replacements have also been done randomly without using a language model (Haque et al., 2021; Shi et al., 2022; Wang et al., 2018). Yet, without context-aware replacement strategies, randomly swapping terms could corrupt the semantics of the text.
Alternatively, some studies duplicate text in a corpus and use a model to paraphrase them (Shi et al., 2022; Guo et al., 2021a; Gao et al., 2020). This could be an effective way of bolstering a dataset for LRLs, as synthesising new sentences that conserve the original meaning of the sentence would expose the language model to more diverse and representative data. A related approach involves adopting grammar correction to improve generated text for question-answering in Tibetan (Sun et al., 2019). This BERT-based strategy for fixing grammatical errors in output sentences could help to reduce erroneous text generation for technical methods that involve randomly replacing terms or performing contextually unaware sentence restructuring. It could be an additional step to ensure that the grammar of generated text is not corrupted. However, it is worth highlighting that this would hinge on the availability and efficacy of grammar correction tools and models. Another study performs corpus expansion by using a reverse transcription model to generate new sentences through grammatical reordering of the original sentences (Wu and Wang, 2023). Given that some languages are more flexible in terms of their word order (Dryer, 2007), enlisting a model to generate grammatically correct alternative sentences could enrich a dataset of LRL text by increasing the diversity of language. Other studies have synthesised data by reordering target language text to align with the structure of a source language (Haque et al., 2021). While this is a beneficial approach for machine translation tasks, this approach could produce structurally incorrect sentences that do not conform to the morphological features of the language. In terms of a dialogue model, training on text data with instances of poor grammar could create a model that is not representative of the language itself. One approach for augmenting and synthesising parallel data involves performing mass translation of text data from a source language to a target language using tools such as GPT 3.5 and Google Translate API (Shafayat et al., 2024).
Another approach uses translation tools to impute missing text examples in parallel datasets (Haque et al., 2021). While these papers use translation tools to enhance translation models, the same approach could be adapted to translate textual data from a higher-resourced language like Chinese to a target LRL. However, the quality of the synthesised data hinges on the existence and availability of performant translation models. Building generative language models on translated text could propagate the errors from translation models through to generative language models. One study references building user-generated text for NMT by employing noise and stylistic changes to textual data (Dandapat and Federmann, 2018). While this method produces synthetic data, it could be more suited for the specific task of translating user-generated content, as opposed to augmenting data for other text generation tasks such as question-answering, whereby noise could distort information.
@@ 4.2.2. Back-Translation
Back–translation is the second most frequently used text data augmentation technique in this review for tackling data scarcity when building generative text models for LRLs (Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Pham et al., 2021; Shi et al., 2022; Marie and Fujita, 2020; Wang et al., 2018; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Wang et al., 2022). While round-trip back-translation can bolster monolingual data (Guo et al., 2021a), back-translation can be a beneficial tool for building synthetic parallel datasets for translation tasks. However, back-translation suffers from similar pitfalls as any other translation-based text data augmentation such as mass translation. The efficacy of back-translation is heavily dependent on the quality and performance of the translation tools that are being used to synthesise these back-translated inputs. While back-translation has improved translation performance for modelling a variety of LRLs (Berckmann and Hiziroglu, 2020; Pham Van and Le Thanh, 2022; Nissanka et al., 2020), this process could consequently produce synthetic text data that disrupts the semantics and grammar of the original text data. Further investigations into boosting LRL data for generative language modelling through the use of back-translation could be an insightful experiment.
@@ 4.2.3. Multilingual Models
Building multilingual models is an approach adopted by many of the papers in this study to address the lack of training data for LRLs. Given that NMT models require a mapping from a source language to a target language and vice versa, there are many NMT models discussed in this review (Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Guo et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Liu, 2022; Yu et al., 2020; Phan-Vu et al., 2017). Cross-lingual transfer learning can be leveraged for various generation tasks by building models for a variety of LRLs (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Downey et al., 2024; Cahyawijaya et al., 2021; Guo et al., 2024b; Sorokin et al., 2022; Gerz et al., 2018; Agarwal et al., 2022; Wongso et al., 2023; Li et al., 2022a). Although one study trains multilingual models from a mixture of 250 high and LRLs (Cahyawijaya et al., 2021), most of the studies create models for related languages (Niyogi and Bhattacharya, 2024; Downey et al., 2024; Cahyawijaya et al., 2021; Wongso et al., 2023; Li et al., 2022a). The study that experiments with 250 high and low resource-languages (Cahyawijaya et al., 2021) argues that building multilingual models can be likened to augmenting the training dataset by 33%. The study explores the potential benefits of training similar or maximally related languages together, suggesting that this approach could enhance performance. These findings present an alternative to monolingual text data augmentation strategies for addressing data scarcity. However, the results also indicate that some multilingual language models can underperform when compared to their monolingual counterparts, highlighting the importance of selecting compatible language groups for multilingual modelling. Alternatively, a pair of studies train models on multiple languages to exploit cross-lingual information (Sorokin et al., 2022; Agarwal et al., 2022). Both of these approaches involve collating information from more resourced languages to generate more relevant and accurate text in a target language. However, this process hinges on such relevant information existing in other languages. Monolingual text data augmentations cannot synthesise new information, therefore this cross-lingual approach could expose LRL models to more diverse information. For example, some LRLs may lack news services that provide coverage of global affairs in their language. Consequently, facilitating cross-lingual information sharing can help bridge this gap, ensuring that speakers of these languages have access to a broader range of knowledge not written explicitly in the target language, such as global news. Another study investigates training a sequence-to-sequence model using parallel data to generate sentences consisting of a mix of both languages (Bhowmick et al., 2023). While this could create seemingly nonsensical data, code-mixing could accurately reflect some communities such as Irish people, as they often inject English into their Irish conversations and vice versa (Laoire, 2016). Therefore, code-mixing could be a valid method of augmenting data for a pair of languages that are frequently used in conjunction with each other.
@@ 4.2.4. Prompt Engineering
Several papers in this study attempt prompt engineering techniques on existing models to investigate their capacity to adapt to different LRLs (Li et al., 2024b; Zhang et al., 2024a; Guo et al., 2024b; Bendel and Jabou, 2024; Agarwal et al., 2023; Saini and Vidhyarthi, 2023; Scalvini and Debess, 2024). An interesting approach enlists a programmer-interpreter process to continuously refine the output of a language model by using smaller models to perform corrections and further changes (Li et al., 2024b). While this process could be beneficial for synthesising additional training data for LRLs with a lack of training data, it can produce nonsensical text that does not conform with the grammatical and semantic features of the target language. Two papers in this review enrich generative language prompts by injecting additional information into the prompt itself (Guo et al., 2024b; Zhang et al., 2024a). One of these approaches involves the Talent framework whereby a textbook is created for an LRL (Guo et al., 2024b). This textbook is subsequently utilised to guide large language models to translate by including vocabulary and language examples with the prompt. While this approach reflects how humans learn a new language, it hinges on the multilingualism of the large language model. Similarly, another study uses a few-shot approach that appends example translations and definitions to guide the translation model (Zhang et al., 2024a). If these models being prompted and guided do not have cross-lingual knowledge or a broad vocabulary range that can represent LRLs, they could struggle to adapt to a language regardless of additional information and context in the prompt. There are several studies in this review that simply assess the capability of existing pre-trained generative language models for a variety of LRLs (Bendel and Jabou, 2024; Agarwal et al., 2023; Saini and Vidhyarthi, 2023; Scalvini and Debess, 2024). A study evaluating GPT-SW3 for Faroese makes the case for using a multilingual language model for a family of languages to augment data for a lower resource language within the family (Scalvini and Debess, 2024). This is an interesting perspective that warrants experimentation. Further research is required to investigate the efficacy of training a language model for an LRL by using such a dataset augmented with text generated by a multilingual model from the same language family.
@@ 4.2.5. Monolingual Models
In contrast with building multilingual models, some studies assess whether existing monolingual data is sufficient for modelling LRLs or if the scarcity of such data prevents effective modelling of LRLs (Acikgoz et al., 2024; Sun et al., 2019; Ulčar and Robnik-Šikonja, 2023; Rizqullah et al., 2023; Salim et al., 2023). An insightful approach involves enlisting transfer learning to adapt existing generative language models for a specific LRL. Using this approach, a Turkish language model was created by exposing existing models to additional training data (Acikgoz et al., 2024). Exploiting transfer learning allowed the model to leverage knowledge from the original languages it was trained on, to better understand and generate text in Turkish, even with limited Turkish-specific data. The paper found that adapting the Mistral model for Turkish improved accuracy across reasoning and question-answering tasks, indicating that adaptive learning can improve LRL modelling. On the other hand, one paper further trains various existing generative models on ChatGPT-generated QA pairs for the Indonesian language (Rizqullah et al., 2023). However, their use of text data sourced from ChatGPT biases their model. Their model is bounded by ChatGPT’s capacity to generate text in the Indonesian language. Similarly, the language used by ChatGPT to create questions might not reflect questions that would be posed by humans. Despite these limitations, using data generated by an existing language model could serve as a foundation and baseline for building generative models for LRLs.
An alternative approach builds a QA corpus generation model for the Tibetan language (Sun et al., 2019). Although it employs BERT as a method to correct grammatical errors in the generated QA text, the paper references that the generation model is not aware of and does not adhere to the grammatical characteristics of the Tibetan language. Their approach randomly masks words in their dataset and uses BERT to replace the term. While this could generate valid grammatical corrections, the model could produce nonsensical or even unrelated terms that lose the context of the original term. These replacements could also corrupt the grammar of the sentences. This word-level approach does not account for wider sentence-level grammatical errors. Another approach involves training a sequence-to-sequence model for the Slovenian language (Ulčar and Robnik-Šikonja, 2023). Interestingly, the paper finds that their small T5 models outperform the multilingual models, yet the large multilingual model outperforms their large T5 model. This highlights that there is some evidence that monolingual models can outperform multilingual models for less-resourced languages. However, these results also suggest that some languages could achieve better performance by leveraging multilingualism and transfer learning. The Bangla language is modelled using a considerable amount of data without using any specific text data augmentation techniques (Salim et al., 2023). However, there are no baseline comparisons in this study. It would be insightful to compare the performance of the monolingual Bangla model with other multilingual models that have been trained on Bangla text.

### v2

@@ 4.1. RQ1: What Low-Resource Languages Have Already Been Modelled by Other Studies?
The findings show a lack of a consistent and universal definition for what constitutes a low-resource language. In terms of low-resource being linked to data availability, some languages have relatively rich digital corpora compared with most underrepresented languages like Aymara. However, the concept of ”low-resource” must also consider the availability of computational infrastructure for modelling underrepresented languages. The findings suggest that pre-training from scratch is rarely explored, with most papers fine-tuning models or prompting existing generative models. This lack of pre-training from scratch could indicate that there is also a technical limitation in researching low-resource languages. Additionally, research interest and the availability of researchers with technical expertise likely contributes to whether a language is low-resource. In other words, a language may remain “low-resource” not only due to data scarcity but also due to the absence of researchers or institutions capable of modelling it. That is to say that countries with stronger academic and technical ecosystems would have more opportunities to model their languages. Therefore, a language should be considered low-resource when it lacks any combination of data, computational power, or researchers with relevant expertise.
The predominance of Indo-European languages in the literature indicates a disparity among the LRLs themselves in terms of their representation. This is further evidenced by the 15 language families (such as the Aymaran and Tupian language families) that were only represented by a single language - versus the 133 Indo-European languages represented in the literature. This distribution could be attributed to region-specific disparities in research capacity. With most of the Indo-European languages belonging to European language branches, this suggests that more resources are available to researchers in Europe. The EU’s commitment to the European Language Equality project and overall research and development (R&D) funding (eg: via Horizon Europe) is indicative of strong support for Indo-European low-resource language technologies (Rehm and Way, 2023). In contrast, the only Nilotic language represented (Chang et al., 2023), South Western Dinka, is spoken by an ethnic group in South Sudan (Lienhardt et al., 1958). To date, no official data has been published detailing South Sudan’s R&D spending as a percentage of GDP (The World Bank, 2024). This highlights the geopolitical and economic factors shaping which languages are prioritised in language modelling.
The lack of research on entire families of LRLs is particularly concerning given that the increasing dominance of AI tools could further increase language inequality in minority and underrepresented communities. With companies and governments alike looking to inject AI technology into their services (Department of Public Expenditure, Infrastructure, Public Service Reform and Digitalisation, 2025; QuantumBlack, AI by McKinsey, 2025), the lack of research for modelling entire families could contribute to technological marginalisation. Given that language models could be productivity-enhancing (Chatterjee et al., 2024), their incompatibility with entire language families could introduce a new dimension of economic inequality. These LRL-agnostic technologies could widen the economic gap between regions that can fully leverage this technology and those that cannot. A lack of basic digital infrastructure like keyboards or unicode support for LRLs compounds the issue of language inequality (Zaugg et al., 2022). This lack of digital infrastructure presents a major challenge in sourcing sufficient data and building LRL models.
Given our findings surrounding the low-resource languages modelled by other papers, we recommend numerous areas for future research. Most notably, a more comprehensive dataset of LRLs should be created that accounts for the availability of data, computational resources and academic structures/technical expertise, in line with our proposed definition of ”low-resource”. This type of research could arm LRL NLP researchers with a strong foundation on the opportunities and challenges facing a given LRL. Furthermore, our review has found that there are benefits to building LRL models for a family of related languages. Tackling LRL modelling from a language family perspective could boost the representation of underrepresented languages. In addition, a language family-based approach would allow communities to pool computational resources and technical expertise. Therefore, further research is required on family-based language modelling as a sustainable and inclusive strategy for addressing language inequality in NLP.
@@
Similarly to the results section for RQ2, the following discussion and meta-analysis is divided by technical methods used to overcome data scarcity in order of frequency, starting with the most commonly used approaches.
@@ 4.2.1. Monolingual Text Data Augmentation
The process of synthesising new data from monolingual text data is the most frequently used approach for overcoming data scarcity. While techniques ranging from simple word replacements to sophisticated paraphrasing and grammatical reordering are used, the effectiveness of these methods varies significantly depending on resource availability and task specificity.
The use of word-level replacements such as synonym and low-frequency term substitutions on existing data is limited by the replacement strategy. Even going beyond simple synonym replacement and using neural approaches like BERT to contextually replace terms can still corrupt the original meaning of the text data or break sentence-level agreement. Improvements such as using POS tags to substitute terms using a paraphrase table could work for some LRLs. Yet, it could be challenging to apply this to LRLs with a complex morphological structure such as Gaeilge, where altering a term in a sentence could corrupt the grammatical agreement of the sentence (Legate, 1999). We propose that performing POS-informed replacements with a contextualised model like BERT could offer a reduction in sentence-level errors for LRLs where specific POS tags do not affect agreement. One of the main flaws of word-level replacements is that this method does not introduce or expose a model to new semantic content. While such methods may increase lexical diversity or syntactic variation, they do not expose the model to new facts, knowledge, or domain-relevant information. As a result, these techniques may be insufficient for tasks that require broader world knowledge or domain-specific understanding, such as QA. Therefore, word-level replacements could suit augmenting training data for translation tasks over other generative tasks like dialogue generation or abstractive summarisation.
Alternatively, sentence-level transformations, often using some form of grammar correction, are also common methods for overcoming data scarcity. Similarly to the benefits of word-level augmentations, sentence-level transformations can increase sentence diversity in the training data. However, sentence-level transformations differ as they aim to preserve the meaning and grammatical features of text. While neural grammar correction strategies have been used to improve the quality of generated text (Sun et al., 2019), the improvement hinges heavily on the availability and efficacy of the grammar correction tools. Nonetheless, effective grammar correction tools could enhance both word and sentence-level transformations. Reducing the likelihood of grammatical agreement errors in generated text via these kinds of grammar tools could make these synthetic sentences more suitable as training data for generative tasks like QA. Alternatively, other papers have experimented with reordering the sentences to match the syntax of target languages. However, using a reverse transcription model to generate new sentences through grammatical reordering of the original sentences could fail for languages with strict word order constraints. Given that some languages are more flexible in terms of their word order (Dryer, 2007), adopting a model to generate grammatically correct alternative sentences could enrich a dataset of LRL text by increasing the diversity of language. This could be useful for languages with dialects that alter the word order as it could produce dialectal variations from the original sentences. Therefore, these grammatical-based sentence-level transformations could be beneficial for translation tasks and could be explored for translations between dialects.
While these substitution methods can benefit LRL modelling, sentence-level transformations with grammar correction capabilities offer the most potential across a wider range of generative tasks. Future work should explore task-aligned augmentation pipelines that go beyond surface-level variation and actively incorporate new semantic information into training data.
@@ 4.2.2. Back-Translation
Given that back-translation is the second most frequently used data augmentation technique and translation is the most common NLG task, this suggests that translation is of particular importance to minority language communities. The widespread use of back-translation underscores the extent of data scarcity and institutional neglect facing LRLs in global language technology development.
Similarly to other translation-based data augmentation techniques, the efficacy of back-translation is heavily dependent on the quality and performance of the translation tools that are being used. While back-translation has improved translation performance for modelling a variety of LRLs (Berckmann and Hiziroglu, 2020; Pham Van and Le Thanh, 2022; Nissanka et al., 2020), this process could consequently produce synthetic text data that disrupts the original semantics and grammar (Park et al., 2024). We propose that sentence-level grammar correction methods could alleviate grammatical errors produced via back-translation. Given that modelling groups of related languages can improve LRL modelling performance (Cahyawijaya et al., 2021), back-translation could also be used to augment data across a variety of LRLs. In particular, this could benefit modelling dialects of LRLs that differ syntactically from the standard form of a language. We propose that translating dialectal Irish to English, and then back to standard Irish, could yield parallel sentence pairs suitable for training dialect-aware generative models.
Overall, back-translation can benefit translation tasks but semantics can be lost in translation, suggesting it would not generalise well for augmenting summarisation datasets where translation errors could contribute to inaccuracies in output. Future research should explore the role of back-translation in augmenting datasets for dialect modelling and language family-level training and investigate hybrid pipelines that combine translation with grammar correction to ensure data quality across tasks.
@@ 4.2.3. Multilingual Models
Building multilingual models is an approach adopted by many of the papers in this study to address the lack of training data for LRLs. However, its more frequent use over monolingual modelling may reflect the need for data pooling across different languages to accurately model LRLs.
For LRLs, the cross-lingual transfer gained by modelling different languages together can improve performance across different tasks (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Downey et al., 2024; Cahyawijaya et al., 2021; Guo et al., 2024b; Sorokin et al., 2022; Gerz et al., 2018; Agarwal et al., 2022; Wongso et al., 2023; Li et al., 2022a). Given that building multilingual models has been likened to augmenting the training dataset by 33% (Cahyawijaya et al., 2021), it could be described as an implicit form of data augmentation. Although synthetic data is not created, this inclusion of multilingual data exposes models to a broader set of language instances. As a result, this suggests more work should be done to pool text data resources between related languages for optimal model training.
However, some evidence shows that some smaller multilingual models have been outperformed by their monolingual counterparts (Ulčar and Robnik-Šikonja, 2023), suggesting that multilingual models perform better as they scale. This tradeoff could highlight the importance of selecting compatible language groups and having sufficient data between the different languages to avoid overfitting. Optimising language groupings remains an open research challenge.
Similarly, the use of code mixed sentences to generate translations could be extended to model languages or dialects of languages that often reflect authentic language practices in bilingual communities. While this could create seemingly nonsensical data, code-mixing could accurately reflect some communities such as Irish people, as they often inject English into their Irish conversations and vice versa (Laoire, 2016). Therefore, code-mixing could be a valid method of augmenting data for a pair of languages that are frequently used in conjunction with each other.
In addition to structural transfer, several studies have used cross-lingual knowledge transfer to exploit information across different language sources (Sorokin et al., 2022; Agarwal et al., 2022). Collating information from higher-resourced languages to generate more relevant text in a target language presents some opportunities and challenges. Previous critiques of monolingual and back-translation augmentations target their inability to synthesise new semantic content. Therefore, this cross-lingual knowledge-sharing approach could expose LRL models to more diverse information that would otherwise not exist in the target language dataset. For example, some LRLs may lack news services that provide coverage of global affairs in their language, preventing this information from existing in a dataset for target LRLs. Consequently, facilitating cross-lingual information sharing could help bridge this gap, as source information could be rephrased in a target LRL. This would ensure that speakers of these languages have access to a broader range of knowledge not written explicitly in the target language.
Future work could involve investigating the use of cross-lingual transfer to build generative models for LRL communities that code-mix between languages. This could be particularly beneficial for LRL languages that are code mixed with more resourced languages like Irish and English, as this could alleviate the data scarcity issue. In addition, more research is required on grouping data resources across related languages so that LRL modelling can benefit from cross-lingual knowledge transfer.
@@ 4.2.4. Prompt Engineering
Prompt engineering has appeared as a method for overcoming data scarcity in several papers. The approach is likely popular given that it is less resource intensive in comparison with model fine-tuning or pre-training from scratch. Its popularity may reflect growing interest in evaluating whether existing generative tools can effectively support LRLs. While prompting existing models might be more accessible, it has its downsides. No matter the prompting approach, generation performance hinges on the multilingualism of the language model being prompted. If a prompted model has been exposed to relatively little data from a language or if the model lacks adequate tokenisation support for a language’s script, then it is likely that the model would struggle to generate accurate text.
Two approaches inject additional information into prompts about the target language and provide examples of translations to guide the generation (Guo et al., 2024b; Zhang et al., 2024a). Although enriching prompts with textbook-like information can improve contextual grounding, the effectiveness of such methods is limited by the model’s tendency to ignore or truncate long or complex prompts. Often, lengthy prompts with too much information can result in prompt sections being forgotten or omitted (Yang et al., 2025). This would be a challenge for the approach of using textbook-like information. Ensuring that a prompted model follows the injected instructions is an open challenge across the field of prompt engineering.
We recommend that future work should evaluate existing language models for their capacity to generate text in LRLs. Prompting existing models could serve as a baseline comparison for future fine-tuned or pre-trained models. Additionally, more work is required across the field of language modelling to develop models and decoding strategies that more reliably attend to and follow prompt instructions, especially in multilingual and low-resource settings.
@@ 4.2.5. Monolingual Models
While monolingual modelling is used by some studies, it is not as popular as multilingual modelling. The difference in popularity could suggest that there is growing interest in multilingual modelling as a method to tackle generative LRL modelling.
This process of training models on monolingual data is often done by exploiting adaptive learning techniques. While one study that fine-tuned Mistral on Turkish data improves accuracy across reasoning and QA tasks (Acikgoz et al., 2024), there are several assumptions required for this method to work. Tokenisation support is essential; without it, a model may fail to represent a language’s syntax or morphology accurately. Additionally, a large language model with billions of parameters would require a significant amount of text data to avoid overfitting during the fine-tuning process. Previously discussed augmentation techniques such as monolingual text data augmentations could help with this challenge by producing synthetic data. Therefore, fine-tuning an existing model could remove the need to pre-train a model from scratch, but the issue of data scarcity ultimately remains an open challenge.
Some evidence shows that small monolingual T5 models can outperform small multilingual models for LRLs (Ulčar and Robnik-Šikonja, 2023). While smaller monolingual models can outperform smaller multilingual ones, multilingual models tend to outperform as model size and available training data increase. This suggests that monolingual models could be suitable for modelling extremely low-resource languages, whereas larger multilingual models could perform better where there is considerable data available for a group of related languages. The stronger performance of multilingual models at scale may explain their growing popularity for LRL modelling. Given that some studies do not report appropriate baseline comparisons (Salim et al., 2023), monolingual models could be a suitable baseline benchmark for LRL feasibility. If a model trained on monolingual data can perform well, the resource gap for the language may not be as large as perceived.
We recommend that future work should focus on using monolingual modelling as a baseline benchmark for whether a language is as low-resourced as perceived. The prevalence of adaptive learning and the data requirements necessitate the collection of more data for LRLs to avoid overfitting. Future research could also investigate the tipping point at which multilingual models begin to consistently outperform monolingual models for LRLs, depending on model size, data availability, and language similarity.